OpenStudy (babyx3boo):

Simplify: 2/3+1/5 over 2/3-1/5

OpenStudy (anonymous):

easiest thing to do here is to multiply top and bottom by 15 to get rid of the compound fraction

OpenStudy (anonymous):

\[\frac{\frac{2}{3}+\frac{1}{5}}{\frac{2}{3}-\frac{1}{5}}\times \frac{15}{15}\]

OpenStudy (anonymous):

then you will have no fractions top and bottom and you will be done once you add

OpenStudy (anonymous):

is it clear where i got the 15 from?

OpenStudy (babyx3boo):

yess is the answer 13/7?

OpenStudy (anonymous):

yeah \[10+3\] up top and \(10-3\) in the bottom

OpenStudy (babyx3boo):

alrighty thank you :)

OpenStudy (anonymous):

yw